OpenStudy (anonymous):

Physics (simple trig): a novice pilot sets a plane's controls, thinking that the plane will fly 250 km/hr to the north. If the wind blows at 75 km/hr to the south east (135 degrees), what is the plane's resultant velocity?

OpenStudy (anonymous):

you add the vectors.  add the components individually. the resulting velocity in the x direction is: 0 + 75cos(-45) = 53.03 in y direction: 250 + 75sin(-45) = 196.97 final = [53.03, 196.97] \[Magnitude = \left| {\sqrt{(53.03)^2 + (196.97)^2}} \right|\] \[Angle = \tan^{-1}\left( \frac{ 196.97 }{ 53.03 } \right)\]